OpenStudy (anonymous):

The adjacent sides of a parallelogram measure 14 centimeters and 20 centimeters, and one angle measures 57 degrees? What is the area of the parallelogram?

OpenStudy (campbell_st):

start by drawing a diagram, or using the diagram  so there are 2 ways to do this problem. 1. find the perpendicular height of the parallelogram \[\sin(57) = \frac{h}{14}\] so \[h = 14 \times \sin(57)\] then use the area formula \[A = base ~~ \times ~~ height\]

OpenStudy (campbell_st):

|a 2nd method involves using the trig formula for area after drawing a diagonal. \[A = \frac{1}{2} a \times b \times \sin(C)\] |dw:1415308882850:dw where a = 14, b = 20 and angle C = 57 when you have the area of the triangle, just double it for the area of the parallelogram. hope it helps

OpenStudy (anonymous):

how would i find the base since you have to make a right triangle? @campbell_st

OpenStudy (campbell_st):

the base remains 20